OpenStudy (anonymous):

Which of the following parent functions has the most inclusive range? A.y=1/x B.y= the square root of x C.y=2^x D.y=x^3

OpenStudy (anonymous):

what does "inclusive range" mean?

OpenStudy (anonymous):

\[y=x^3\] has range \((-\infty, \infty)\) i.e \(\mathbb{R}\) but that doesn't mean the answer is right, since i have never heard the term "inclusive range" before it does have the biggest one however